# appReferences Trajectory-aware Shifted State Space Models for Online Video Super-Resolution

### v1

@@ Abstract
Online video super-resolution (VSR) is an important technique for many real-world video processing applications, which aims to restore the current high-resolution video frame based on temporally previous frames. Most of the existing online VSR methods solely employ one neighboring previous frame to achieve temporal alignment, which limits long-range temporal modeling of videos. Recently, state space models (SSMs) have been proposed with linear computational complexity and a global receptive field, which significantly improve computational efficiency and performance. In this context, this paper presents a novel online VSR method based on Trajectory-aware Shifted SSMs (TS-Mamba), leveraging both long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. Specifically, TS-Mamba first constructs the trajectories within a video to select the most similar tokens from the previous frames. Then, a Trajectory-aware Shifted Mamba Aggregation (TSMA) module consisting of proposed shifted SSMs blocks is employed to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and corresponding shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Additionally, we propose a trajectory-aware loss function to supervise the trajectory generation, ensuring the accuracy of token selection when training our model. Extensive experiments on three widely used VSR test datasets demonstrate that compared with six online VSR benchmark models, our TS-Mamba achieves state-of-the-art performance in most cases and over 22.7% complexity reduction (in MACs). The source code for TS-Mamba will be available at https://github.com.
Introduction
@@
Among various video super-resolution (VSR) application scenarios, online VSR has recently attracted significant interest due to the growing popularity of live video conferencing and live broadcasting applications (Fuoli et al. 2023; Xiao et al. 2023). In online VSR, the current high-resolution (HR) video frame is typically restored using only its low-resolution (LR) counterpart and previous frames. This is constrained by the requirements for low latency and low computational complexity inherent to these online real-time applications.
@@
In a VSR model, temporal alignment or aggregation is a core module employed to compensate for the information from neighboring frames before generating the current HR frame. Advanced temporal alignment or aggregation modules have been recently developed, which are based on deformable convolution networks (DCN) (Wang et al. 2019; Tian et al. 2020), flow-guided deformable alignment models (Chan et al. 2022; Zhu et al. 2024b), non-local attention mechanisms (Li et al. 2020; Yi et al. 2019) or Vision Transformer based spatio-temporal information aggregation (Liu et al. 2022a; Tang et al. 2023). Although they offer superior VSR performance, these methods are typically associated with high complexity and, therefore, are not ideal for online VSR.
@@
To mitigate these limitations, recent online VSR methods have adopted more efficient temporal alignment modules, such as lightweight optical flow networks (Sajjadi, Vemulapalli, and Brown 2018; Xiao et al. 2023), deformable attention mechanisms (Fuoli et al. 2023; Yang et al. 2023), and temporal motion propagation modules (Zhang et al. 2024b). Despite their efficiency, these methods predominantly use short-term temporal information based on convolutional neural networks (CNN) — typically from a single previous frame, which restricts their ability to further enhance reconstruction quality. While incorporating long-term temporal alignment can improve performance, it often introduces significant computational overhead, resulting in challenges for real-time or resource-constrained applications.
@@
Recently, low-complexity state space models (SSMs) (Gu, Goel, and Ré 2021; Gu and Dao 2023) have been proposed with linear computational complexity and with relatively large receptive fields, which can potentially improve performance with limited complexity. Inspired by this, we propose a Trajectory-aware Shifted Mamba for online VSR, denoted as TS-Mamba, leveraging long-term trajectory modeling and low-complexity Mamba for achieving the token-level spatio-temporal aggregation. In TS-Mamba, trajectories within a video are first constructed for selecting the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation (TSMA) module is then employed, which consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and four different shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Moreover, we propose a trajectory-aware loss function to supervise the trajectory generation, optimizing the accuracy of token selection when training our model. The proposed TS-Mamba model enables efficient long-term video modeling with significantly reduced computational complexity. The primary contributions are summarized as follows:
@@
The novel shifted SSMs blocks are designed based on four different shift operations and Hilbert scannings to effectively compensate for the intra-window and inter-window losses of Hilbert scannings and strengthen local spatial continuity of Mamba.
@@
Related Work
Video Super-Resolution
@@
Video super-resolution (VSR) is a fundamental low-level vision task that aims to restore an HR video from its LR counterpart. Existing VSR methods are typically learning-based, utilizing various deep neural networks (Teed and Deng 2020; Zhu et al. 2019; Arnab et al. 2021; Ho et al. 2022). For example, optical flow-based methods (Chan et al. 2021; Liu et al. 2022b) explore the temporal motion between frames to align them; deformable convolution networks (DCN)-based methods (Tian et al. 2020; Wang et al. 2019) learn the motion offsets between frames for feature alignment. Moreover, flow-guided deformable-based methods (Chan et al. 2022; Zhu et al. 2024b) combine optical flow and DCN to achieve better feature alignment. Non-local attention-based methods (Li et al. 2020; Yi et al. 2019) aggregate global information for feature aggregation. Vision Transformer-based methods (Liu et al. 2022a; Tang et al. 2023) aggregate long-term spatio-temporal information in video to restore SR frames. However, these methods are often associated with high complexity and are therefore not best suited for online VSR.
Online Video Super-Resolution
@@
Due to the specific requirement of online applications, online VSR methods are expected to be lightweight and have low latency. Therefore, most existing online VSR methods have been proposed (Fuoli et al. 2023; Sajjadi, Vemulapalli, and Brown 2018; Xiao et al. 2023) with efficient feature alignment modules. For example, DAP (Fuoli et al. 2023) designed a deformable attention pyramid module to dynamically focus on the most salient locations between frames. FRVSR (Sajjadi, Vemulapalli, and Brown 2018) and CKBG (Xiao et al. 2023) utilized lightweight optical flow networks to estimate motion between frames and perform motion compensation. KSNet (Jin et al. 2023) proposed a kernel-split manner to reparameterize convolutional kernels on the high-value channel, enabling representation of dynamic information and reducing complexity along the channel dimension. FDAN (Yang et al. 2023) proposed a flow-guided deformable attention propagation module to efficiently exploit the temporal information between frames. TMP (Zhang et al. 2024b) employs an efficient temporal motion propagation method that leverages motion field continuity to achieve fast feature alignment. It is noted that, however, these online VSR methods are only based on one previous frame in feature alignment due to the complexity limitation, which hinders further improvement of VSR performance.
State Space Models
@@
State space models (Gu, Goel, and Ré 2021; Gu and Dao 2023), e.g., Mamba, have been widely employed in vision tasks (Liu et al. 2024; Zhu et al. 2024a) due to their linear computational complexity and ability to model global dependencies. Mamba typically converts 2D images into 1D tokens through scanning (Qiao et al. 2024; Shi et al. 2025), resulting in spatial continuity loss inherent to images. More recently, advanced scanning techniques have emerged to address this issue, such as bidirectional scanning (Hu et al. 2024; Shi et al. 2025), cross scanning (Liu et al. 2024), continuous 2D scanning (Yang et al. 2024), and local scanning (Huang et al. 2024). To the best of our knowledge, the use of Mamba has not yet been investigated for the online video super-resolution task. Unlike existing Mamba-based works, we introduce sophisticated shift operations for Hilbert scannings to enhance the ability of Mamba to maintain local spatial continuity.
Methodology
@@
where consists of a convolution layer and residual blocks to generate features and tokens from video frames, is the token number, and is the temporal window size.
Based on the generated tokens , the trajectories of frame can be formulated as a set of trajectories,
@@
Finally, the aggregated feature and the current LR frame are fed into the reconstruction network and the upsampling network, respectively, to produce the super-resolved frame :
@@
Token Selection
In order to select most similar tokens along trajectories, we first reformulate tokens , associated with trajectories . Based on the formulation of the trajectories in Equation (3), tokens , and can be formulated as:
We compute the cosine similarity between the token , and tokens to select the most similar tokens along trajectories. The indices of the selected tokens and the selected tokens can be formulated as:
@@
Trajectory-aware Shifted Mamba Aggregation
Mamba networks are typically used to convert 2D images into 1D tokens via scanning, resulting in spatial continuity losses inherent to the images. Existing works (Zhang et al. 2024a; Xiao and Wang 2025) do not analyze the degree of discontinuous regions but instead repeatedly use multiple scannings, making these methods hard to maintain the spatial continuity of the image and instead lead to greater complexity.
@@
Discontinuity for Hilbert Scanning
To evaluate the spatial discontinuity of Hilbert scannings in local windows, we define the discontinuity degree as follows. If the four adjacent areas are successively scanned, the region consisting of these four scanned areas is considered as continuous ( = 0); otherwise, the discontinuity degree equals the number of areas that are not successively scanned. For a region consisting of four adjacent areas, the range of discontinuity degree is . This is illustrated in Figure 3 (a), where four typical Hilbert scannings, i.e., Scan-1, Scan-2, Scan-3 and Scan-4 are shown on a 44 grid. Here the region with = 1 is marked by a green circle and the region with = 2 is marked by a red circle.
@@
Shifted SSMs Block
To eliminate the discontinuity of Hilbert scannings, we propose the “Scan-Shift-Scan” manner that combines window shifting with specific Hilbert scannings to strengthen the continuity of SSMs. The shifting can be defined based on the shift direction and shift position, e.g., Up 1 position ((1)), Up Left 1 position ((1)) and Down Right 2 position ((2)). Our “Scan-Shift-Scan” manner is designed based on the four Hilbert scannings (shown in Figure 3 (a)) and these window shifting processes. As shown in Figure 3 (b), we illustrate the procedure of Scan-1Scan-3 as an example. The local windows are first partitioned by shift operation and then cyclic fed as the shifted windows. It can be inferred that the second scanning (Scan-3) on the shifted window can eliminate the discontinuity of first scanning (Scan-1).
To evaluate the discontinuity elimination, we set three symbols and define an elimination value to mark and calculate the elimination. Specifically, we use green“\”, red“\”, and gray“\” on the circle for representation that eliminates 1, 2, and 3 discontinuity degrees, respectively, in Figure 3 (b)-(c). The elimination value is calculated by summing the eliminated discontinuity degrees that consist of intra-window discontinuity elimination and inter-window discontinuity elimination, i.e., . We have investigated many combinations of shift operations and scannings, and illustrate the representative shifted windows generated by six shift operations, i.e., (1), (2), (3), and (1), (2), (3), under the first scanning (Scan-1) and second scanning (Scan-3) in Figure 3 (c). It can be inferred from Figure 3 (b)-(c) that the procedure of Scan-1Scan-3 achieves the best elimination (=18) under the first scanning is Scan-1. This procedure achieves the best intra-window discontinuity elimination (=18) but doesn’t eliminate inter-window discontinuity (=0). We can infer that the other three procedures can also achieve the best elimination: Scan-2Scan-4, Scan-3 Scan-1, Scan-4Scan-2. Moreover, the procedure of Scan-1Scan-3 has the best inter-window discontinuity elimination (=6) but worse than the procedure of Scan-1Scan-3 for intra-window discontinuity elimination (=8). and shift operations exhibit symmetry under the first scanning Scan-1 and the second scanning Scan-3 when the same shift positions. Different combinations of shift operations and scannings can bring significant different elimination performance (more detail in our supplementary).

### v2

# Trajectory-aware Shifted State Space Models for Online Video Super-Resolution

@@ Abstract
Online video super-resolution (VSR) is an important technique for many real-world video processing applications, which aims to restore the current high-resolution video frame based on temporally previous frames. Most of the existing online VSR methods solely employ one neighboring previous frame to achieve temporal alignment, which limits long-range temporal modeling of videos. Recently, state space models (SSMs) have been proposed with linear computational complexity and a global receptive field, which significantly improve computational efficiency and performance. In this context, this paper presents a novel online VSR method based on Trajectory-aware Shifted SSMs (TS-Mamba), leveraging both long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. Specifically, TS-Mamba first constructs the trajectories within a video to select the most similar tokens from the previous frames. Then, a Trajectory-aware Shifted Mamba Aggregation (TSMA) module consisting of proposed shifted SSMs blocks is employed to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and corresponding shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Additionally, we propose a trajectory-aware loss function to supervise the trajectory generation, ensuring the accuracy of token selection when training our model. Extensive experiments on three widely used VSR test datasets demonstrate that compared with six online VSR benchmark models, our TS-Mamba achieves state-of-the-art performance in most cases and over 22.7% complexity reduction (in MACs). The source code for TS-Mamba will be available at https://github.com/QZ1-boy/TS-Mamba.
1 Introduction
@@
Among various video super-resolution (VSR) application scenarios, online VSR has recently attracted significant interest due to the growing popularity of live video conferencing and live broadcasting applications (Fuoli et al., 2023; Xiao et al., 2023). In online VSR, the current high-resolution (HR) video frame is typically restored using only its low-resolution (LR) counterpart and previous frames. This is constrained by the requirements for low latency and low computational complexity inherent to these online real-time applications.
@@
In a VSR model, temporal alignment or aggregation is a core module employed to compensate for the information from neighboring frames before generating the current HR frame. Advanced temporal alignment or aggregation modules have been recently developed, which are based on deformable convolution networks (DCN) (Wang et al., 2019; Tian et al., 2020), flow-guided deformable alignment models (Chan et al., 2022a; Zhu et al., 2024c), non-local attention mechanisms (Li et al., 2020; Yi et al., 2019), Vision Transformer based spatio-temporal information aggregation (Liu et al., 2022a; Tang et al., 2023) or Diffusion models (Wang et al., 2025; Liu et al., 2025; Zhuang et al., 2025). Although they offer superior VSR performance, these methods are typically associated with high complexity and, therefore, are not ideal for online VSR.
@@
To mitigate these limitations, recent online VSR methods have adopted more efficient temporal alignment modules, such as lightweight optical flow networks (Sajjadi et al., 2018; Xiao et al., 2023), deformable attention mechanisms (Fuoli et al., 2023; Yang et al., 2023), and temporal motion propagation modules (Zhang et al., 2024b). For example, CKBG (Xiao et al., 2023) utilized a lightweight optical flow network to estimate motion between frames and perform motion compensation. DAP (Fuoli et al., 2023) designed a deformable attention pyramid module to dynamically focus on the most salient locations between frames and progressive refine the offsets to improve temporal alignment performance. FDAN (Yang et al., 2023) proposed a flow-guided deformable attention propagation module that introduces the optical flow to guide the offset generation to efficiently exploit the temporal information between frames. Despite their efficiency, these methods predominantly use short-term temporal information based on convolutional neural networks (CNN) — typically from a single previous frame, which restricts their ability to further enhance reconstruction quality. While incorporating long-term temporal alignment can improve performance, it often introduces significant computational overhead, resulting in challenges for real-time or resource-constrained applications. Therefore, it is valuable to develop the efficient long-range models for online VSR applications.
@@
Recently, low-complexity state space models (SSMs) (Gu et al., 2021; Gu & Dao, 2023) have been proposed with linear computational complexity and with relatively large receptive fields, which can potentially improve performance with limited complexity. Inspired by this, we propose a Trajectory-aware Shifted Mamba for online VSR, denoted as TS-Mamba, leveraging long-term trajectory modeling and low-complexity Mamba for achieving the token-level spatio-temporal aggregation. In TS-Mamba, trajectories within a video are first constructed for selecting the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation (TSMA) module is then employed that consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed with specific procedures based on Hilbert scannings and four shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Moreover, we propose a trajectory-aware loss function to supervise the trajectory generation, optimizing the accuracy of token selection when training our model. The proposed TS-Mamba model enables efficient long-term video modeling with significantly reduced computational complexity. The primary contributions are summarized as follows:
@@
The novel shifted SSMs blocks are designed based on four different shift operations and Hilbert scannings to effectively compensate for the intra-window and inter-window losses of Hilbert scannings and strengthen the local spatial continuity of Mamba.
@@
2 Related Work
2.1 Video Super-Resolution
@@
Video super-resolution (VSR) is a fundamental low-level vision task that aims to restore an HR video from its LR counterpart. Existing VSR methods are typically learning-based, utilizing various deep neural networks (Teed & Deng, 2020; Zhu et al., 2019; Arnab et al., 2021; Ho et al., 2022). For example, optical flow-based methods (Chan et al., 2021; Liu et al., 2022b) explore the temporal motion between frames to align them; deformable convolution networks (DCN)-based methods (Tian et al., 2020; Wang et al., 2019; Dong et al., 2023) learn the motion offsets between frames for feature alignment. Moreover, flow-guided deformable-based methods (Chan et al., 2022a; Zhu et al., 2024c) combine optical flow and DCN to achieve better feature alignment. Non-local attention-based methods (Li et al., 2020; Yi et al., 2019) aggregate global information for feature aggregation. Vision Transformer-based methods (Liu et al., 2022a; Tang et al., 2023; Lin et al., 2022) aggregate long-term spatio-temporal information in video to restore SR frames. Diffusion-based methods (Wang et al., 2025; Liu et al., 2025; Zhuang et al., 2025) build the long-range modeling using designed diffusion models for information aggregation. However, these methods are still associated with high complexity and are therefore not best suited for real-time online VSR applications. Recently, some Mamba-based works (Xiao & Wang, 2025; Tran et al., 2025) were proposed that use low-complexity Mamba with global receptive filed to improve VSR performance. Although Mamba-based methods achieve some reduction in model complexity, their high-overhead repeated scannings hinders efficient implementation of real-time online VSR.
2.2 Online Video Super-Resolution
@@
Due to the specific requirements of online applications, online VSR methods are expected to be lightweight and have low latency. Therefore, most existing online VSR methods have been proposed (Sajjadi et al., 2018; Cao et al., 2021; Fuoli et al., 2023; Xiao et al., 2023; Jiang et al., 2025a) with efficient feature alignment modules. For example, EGVSR (Cao et al., 2021) and CKBG (Xiao et al., 2023) utilized lightweight optical flow networks to estimate motion between frames and perform motion compensation. KSNet (Jin et al., 2023) proposed a kernel-split manner to reparameterize convolutional kernels on the high-value channel, enabling representation of dynamic information and reducing complexity along the channel dimension. TMP (Zhang et al., 2024b) employs an efficient temporal motion propagation method that leverages motion field continuity to achieve fast feature alignment. DAP (Fuoli et al., 2023) designed a deformable attention pyramid module to dynamically focus on the most salient locations between frames and progressive refine the offsets to achieve temporal alignment improvement. FDAN (Yang et al., 2023) proposed a flow-guided deformable attention propagation module that introduces the optical flow to guide the offset generation to efficiently exploit the temporal information between frames. It is noted that, however, these online VSR methods are only based on one previous frame in feature alignment due to the complexity limitation, which hinders further improvement of VSR performance. Different from existing deformable-based methods (Fuoli et al., 2023; Yang et al., 2023) that based on short-term spatio-temporal aggregation, in this work, our TS-Mamba introduces of trajectories and designs shifted SSMs blocks, enabling it improve the ability for long-range spatio-temporal information aggregation.
2.3 State Space Models
@@
State space models (Gu et al., 2021; Gu & Dao, 2023), e.g., Mamba, have been widely employed in vision tasks (Liu et al., 2024; Zhu et al., 2024a) due to their linear computational complexity and ability to model global dependencies. Recently, some Mamba-based methods are proposed for image/video super-resolution. For example, MambaIR (Guo et al., 2024), and MambaIRv2 (Guo et al., 2025) used Mamba to achieve the global receptive field. TAMambaIR (Peng et al., 2025) introduced a texture-aware state space model to focus on textures regions for improving SR quality. VSRM (Tran et al., 2025) proposed the spatial-to-temporal Mamba and the temporal-to-spatial Mamba to ability of spatio-temporal aggregation. MamEVSR (Xiao & Wang, 2025) proposed a interleaved Mamba and a cross modality Mamba to interleave tokens and further leverage spatio-temporal information to capture finer details. Typically, Mamba converts 2D images into 1D tokens through scanning (Qiao et al., 2024; Shi et al., 2025), resulting in spatial continuity loss inherent to images. To enhance the ability of Mamba, some advanced scanning techniques have emerged to address this issue, such as bidirectional scanning (Hu et al., 2024; Shi et al., 2025), cross scanning (Liu et al., 2024), continuous 2D scanning (Yang et al., 2024a), and local scanning (Huang et al., 2024). To the best of our knowledge, the use of Mamba has not yet been investigated for the online video super-resolution task. Unlike existing Mamba-based works (Xiao & Wang, 2025; Tran et al., 2025) that neglect the local spatial continuity of Mamba, we introduce sophisticated shift operations for Hilbert scannings to enhance the ability of Mamba to maintain local spatial continuity, improving the online VSR performance with high efficiency.
3 Methodology
@@
where consists of a convolution layer and residual blocks to generate features and tokens from video frames, is the token number, and is the temporal window size. Based on the generated tokens , the trajectories of frame can be formulated as a set of trajectories,
@@
Finally, the aggregated feature and current LR frame are fed into the reconstruction network and the upsampling network, respectively, to produce the super-resolved frame :
@@
3.1 Token Selection
In order to select the most similar tokens along trajectories, we first reformulate tokens , associated with trajectories . Based on the formulation of the trajectories in Equation 3, tokens , and can be formulated as:
We compute cosine similarity between the token , and tokens to select the most similar tokens along trajectories. The indices of the selected tokens and the selected tokens can be formulated as:
@@ 3.1 Token Selection
3.2 Trajectory-aware Shifted Mamba Aggregation
Mamba networks are typically used to convert 2D images into 1D tokens via scanning, resulting in spatial continuity losses inherent to the images. Existing works (Zhang et al., 2024a; Xiao & Wang, 2025) do not analyze the degree of discontinuous regions but instead repeatedly use multiple scannings, making these methods hard to maintain the spatial continuity of the image and instead lead to greater complexity.
@@ 3.2 Trajectory-aware Shifted Mamba Aggregation
3.3 Discontinuity for Hilbert Scanning
To evaluate the spatial discontinuity of Hilbert scannings in local windows, we define the discontinuity degree as follows. If the four adjacent areas are successively scanned, the region consisting of these four scanned areas is considered as continuous ( = 0); otherwise, the discontinuity degree equals the number of areas that are not successively scanned. For a region consisting of four adjacent areas, the range of discontinuity degree is . This is illustrated in Figure 3 (a), where four typical Hilbert scannings, i.e., , Scan-2, Scan-3, Scan-4}, are shown on a 44 grid. Here the region with = 1 is marked by a green circle and the region with = 2 is marked by a red circle.
@@ 3.3 Discontinuity for Hilbert Scanning
3.4 Shifted SSMs Block
To eliminate the discontinuity of Hilbert scannings, we propose the “Scan-Shift-Scan” manner that combines window shifting with specific Hilbert scannings to strengthen the continuity of SSMs. The shifting can be defined based on the shift direction and shift position, e.g., Up 1 position ((1)), Up Left 1 position ((1)) and Down Right 2 position ((2)). Our “Scan-Shift-Scan” manner is designed based on the four Hilbert scannings (shown in Figure 3 (a)) and these window shifting processes. As shown in Figure 3 (b), we illustrate the procedure of Scan-1Scan-3 as an example. The local windows are first partitioned by shift operation and then cyclic fed as the shifted windows. It can be inferred that the second scanning (Scan-3) on the shifted window can eliminate the discontinuity of first scanning (Scan-1). To simplify the description of procedure, we define the procedure as:
where the first and the second scannings are denoted as , . Shift operations are denoted as . Therefore, procedure Scan-1Scan-3 can be formulated as .
To evaluate the discontinuity elimination, we set three symbols and define an elimination value to mark and calculate the elimination. Specifically, we use “\”, “\”, and “\” on the circle for representation that eliminates 1, 2, and 3 discontinuity degrees, respectively, in Figure 3 (b)-(c). The elimination value is calculated by summing the eliminated discontinuity degrees that consist of intra-window discontinuity elimination and inter-window discontinuity elimination, i.e., . We have investigated possible procedures and illustrated the representative shifted windows generated by six shift operations, i.e., (1), (2), (3), and (1), (2), (3), under the first scanning (Scan-1) and second scanning (Scan-3) in Figure 3 (c). It can be inferred from Figure 3 (b)-(c) that procedure achieves the best elimination (=18), and the best intra-window discontinuity elimination (=18) but doesn’t eliminate inter-window discontinuity (=0). We infer that the other three procedures also achieve the best elimination: , , . Moreover, procedures , and have the best inter-window discontinuity elimination (=6) but worse than procedure for intra-window discontinuity elimination (=8). The more procedures and details are provided in the supplementary.